# Benchmarking Retrieval-Augmented Large Language Models in Biomedical NLP: Application, Robustness, and Self-Awareness

### v1

@@ 3.1 Settings and Dataset
We evaluated three state-of-the-art LLMs: LLamA2-13B (Touvron et al., 2023) , MedLLamA-13B (Wu et al., 2023), and LLaMA3 8B [^3] https://github.com/meta-llama/llama3, along with three retrievers: BM25 (Luo et al., 2023), Contriver (Izacard et al., 2021), and MedCPT (Su et al., 2023). We considered five biomedical NLP tasks: triple extraction, link prediction, text classification, question answering, and natural language inference, across nine datasets: ADE, ChemProt, GIT, PHarmKG, Hetionet, Ade-corpus-v2, SemedCLass, MedMCQA, and BioNLI. The data statistics are shown in Table 1. The experiments were conducted using A100 GPUs.

### v2

@@ 3.1 Settings and Dataset
We evaluated three state-of-the-art LLMs: LLamA2-13B (Touvron et al., 2023) , MedLLamA-13B (Wu et al., 2023), and LLaMA3 8B [^3] https://github.com/meta-llama/llama3, along with three retrievers: BM25 (Luo et al., 2023), Contriver (Izacard et al., 2021), and MedCPT (Jin et al., 2023). We considered five biomedical NLP tasks: triple extraction, link prediction, text classification, question answering, and natural language inference, across nine datasets: ADE, ChemProt, GIT, PHarmKG, Hetionet, Ade-corpus-v2, SemedCLass, MedMCQA, and BioNLI. The data statistics are shown in Table 1. The experiments were conducted using A100 GPUs.